OpenStudy (anonymous):

Two congruent "door-stop" trapezoids are placed edge-to-edge with measurements and markings as shown. Find the length of

OpenStudy (anonymous):

find the length of 

OpenStudy (anonymous):

The tick in the middle of the segments means their congruent so you just need to find the height of the trapezoid. If you notice, the height along with segment EF form a right triangle with hypotonuse 25, so you can use pythagorean theorem and get \[x^2 + 20^2 = 25^2\] \[x^2 +400 = 625 \] \[x = \sqrt{225}\] \[x = \sqrt{25\times5}\] \[x=5\sqrt{5}\] since its asking your for A and C, you have to double it and get \[AC = 10\sqrt{5}\]

OpenStudy (anonymous):

thank u so much u r great help